Andy (Coordinator): Andy reported that Paul Young has accepted the position of Library Supervisor II at Kilmer Library and will no longer be a member of the CMG. Paul now manages RDS duties and is the night supervisor at Kilmer. He was congratulated for his work in Government Documents at Alexander and LSM and his work with the CMG.
Andy has received a draft of the "marked item use" project overview from Rob for review. A revised draft will be sent to the group for comment in March.
Water emergency supplies for 2008 have been shipped to all units.
Andy reported on the 2008 aged in process report. 1,137 items were searched in our collections and 74 were located. This project may change next year, as the processing for “In-Process” items has changed from an edit to a charge to a user.
Andy informed the group that a link to the Library of Congress Preservation page has been posted to the Collection Management section of Staff Resources (http://www.loc.gov/preserv/).
Elsa (Dana): Collection management staff at Dana are continuing shifts of the Stacks Collection and Periodical Collection and training new students for collection management work.
Mary (Alexander): Mary reported that collection management staff at Alexander has taken on the responsibility of Government Documents work after Paul Young's promotion. Shifts are in the planning stages for Alexander's Government Documents due to the lack of space in certain areas. Mary also reported that they were continuing to shelf read the collections at the Alexander Library.
Kim (Kilmer): Kim reported that there was a water leak in a stairwell, away from the collection. The elevators were out of order recently, so books were hand carried from the circulation desk to the collections on the 2nd floor. Kim also reported on her work with an internal relocate of items from the Reference to the Stacks collection. Kim is working on preparing the Stacks Collection for a large transfer of business titles from the Alexander Library this year.
Rob (Alexander): Rob discussed the two recent water emergencies at Alexander in further detail. In the first water related emergency, a pipe in the East Asian library broke, but no books were in the near area. In the second water related emergency, the third floor Stacks Collection was impacted by a leak over the Winter Holiday. A cleaning project is now underway in this area to clean the shelves and identify moldy items. Rob discussed methods of cutting plastic tarps so that they can be deployed on the collection quickly. He also reported on the shelving status in Government Documents, primarily micro fiche. Rob discussed the need for a collection shift in periodical sections of the collection, Doc NJ and Doc US Law.
Bill (Douglass): Bill currently assisting with collection management responsibilities at the Chang Library during the evenings.
Brian (Alexander): Brian discussed the testing of sprinklers in the building had resulted in some small water leaks.
Eddie (LSM): Eddie reported that Facilities were in the midst of the painting the walls at LSM this past month. The painting resulted in taking down the Periodical sorting shelves for about a week. Kevin Conover has taken over the duties of shelving Government Documents from Paul Young. A J through Z collection mini-shift of the periodical collection is currently underway.
Andy distributed a list of 932 serial titles that have been canceled in print for RUL sites. Collection management staff should use this list to reallocate the available space for growth assigned to these now canceled titles to other active titles nearby. DTS staff will be forwarding this list of canceled titles to the CMG once a year.
Andy distributed a list of 30 titles to group members detailing periodical titles that will use "superseded procedures" to share the holdings for a single title with multiple RUL sites. Titles will be added to this list as decisions are made to share the holdings. CMG staff at non-DTS hub sites will be asked to pull holdings according to directions on a processing slip supplied by DTS. Once DTS has finalized the procedures for these titles, Yuhwei Ling will visit the group to discuss the necessary work.
Rob requested the help of group members in organizing the surplus shelving kept for NBL sites in the BB level of Alexander Library. Surplus shelving that will not be used will be discarded, while newer surplus shelving will be rearranged. This work is tentatively scheduled for Spring Break/2009. An email will go out to group members as we get closer to this date.
Wednesday, March 11, 2009
2:30 – 4:30 pm